# Reinforcement Learning with Backtracking Feedback

### v1

@@ 5 Experimental Results
Table 1: Comparison of IT, RL, BSAFE+ and RLBF models on LMSYS with Middle Filling attacks and a subset of LMSYS with harmful queries.
@@
In this section, we present empirical evidence validating the effectiveness of RL with Backtracking Feedback (RLBF). We conduct a comparative analysis against relevant baselines, including standard Instruction Tuned models (IT), IT models trained on our reward function that excludes backtracking specific rewards, BSAFE+, and Circuit Breakers (Zou et al., 2024), focusing on robustness against adversarial attacks and the preservation of model utility. We provide all necessary information to reproduce the experiments in the supplementary material.
@@ 5.1 Robustness Against Harmful Content Generation
The high ASRs exhibited by the baseline IT models (68%–81% on LMSYS-MF, 24%–28% on LMSYS) underscore the known limitations of standard instruction tuning for robust safety. These models often develop “shallow safety,” easily bypassed by attacks like MF that inject malicious instructions after an initially benign context. The marginal improvements observed with RL (61%–72% on LMSYS-MF, 22%–25% on LMSYS) suggest that conventional RLHF/RLAIF, while potentially reducing direct refusals on benign prompts, does not inherently equip models to handle sophisticated, in-context safety violations without specific mechanisms.
@@
Table 2 extends this analysis to other adversarial strategies: Greedy Coordinate Gradient (GCG) attacks and manipulation of Decoding Parameters. These attacks represent different threat vectors, testing the model’s internal robustness and sensitivity to generation configurations. Against GCG attacks, RLBF consistently achieves the lowest ASR (4.3%–4.7%) compared to all baselines, including the strong Circuit Breakers (10.7%–13.4%) and BSAFE+ (5.7%–6.6%). Similarly, against Decoding Parameter attacks, while both BSAFE+ and RLBF perform exceptionally well (e.g., 1.0% ASR on MaliciousInstruct), RLBF shows a slight edge on the HEx-PHI benchmark (3.7% vs 5.0%). This superior performance against diverse, adaptive attacks further reinforces the benefits of the integrated RL optimization within RLBF, which likely fosters a more fundamental robustness to safety violations beyond what SFT-based correction or external filters alone can achieve.
@@ A.2 Critic Prompt
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification:
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification:
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: No theory is provided in the paper.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: All information necessary to reproduce experiments are supplied in the supplemental.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: At this time, we provided all the needed information to reproduce the results given in the paper. We will consider releasing code upon acceptance.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We supply all the experimental details in the supplemental.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: It is compute infeasible to train multiple version of all our models to provide error bars.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: All sufficient information on the computer resources needed to reproduce the experiments are supplied in the supplemental.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification:
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We provide it in the limitations section.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We have not released new datasets or models at this time.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We provided all the license and terms of use explicitly in the supplemental.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: We have not released new datasets or code at this time.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v2

@@ 4.2.1 LLM Safety Critic
To validate the critic, we manually constructed 100 prompt–response pairs spanning a range of violation locations and severities (43 high, 38 medium, 19 low). Three candidate critic prompts achieved 97.6%, 97.6%, and 98.0% accuracy on this set (three samples per pair), indicating low sensitivity to prompt wording. We examine the effect of critic quality on the trained policy in Section 5.5.
@@ 5 Experimental Results
Table 1: Comparison of IT, RL, Circuit Breakers (CB), BSAFE+ and RLBF models on LMSYS with Middle Filling attacks and a subset of LMSYS with harmful queries.
@@
In this section, we present empirical evidence validating the effectiveness of RL with Backtracking Feedback (RLBF). We conduct a comparative analysis against relevant baselines, including standard Instruction Tuned models (IT), IT models trained on our reward function that excludes backtracking specific rewards, BSAFE+, and Circuit Breakers (Zou et al., 2024), focusing on robustness against adversarial attacks and the preservation of model utility. All experiments use A100 80GB GPUs with bf16 mixed precision and the AdamW optimizer, with a 100-step linear warmup followed by cosine decay. The learning rate is $1\mathrm{e}{-5}$ for SFT and $5\mathrm{e}{-6}$ for RL, and we set $\lambda_{SFT}=0.1$. The SFT dataset contains 300k augmented examples generated as in Section 4.1.2. Training each model required approximately 800 GPU hours (250 for SFT and 550 for RL).
@@ 5.1 Robustness Against Harmful Content Generation
The high ASRs exhibited by the baseline IT models (68%–81% on LMSYS-MF, 24%–28% on LMSYS) underscore the known limitations of standard instruction tuning for robust safety. These models often develop “shallow safety,” easily bypassed by attacks like MF that inject malicious instructions after an initially benign context. The marginal improvements observed with RL (61%–72% on LMSYS-MF, 22%–25% on LMSYS) suggest that conventional RLHF/RLAIF, while potentially reducing direct refusals on benign prompts, does not inherently equip models to handle sophisticated, in-context safety violations without specific mechanisms. Circuit Breakers (13%–18% on LMSYS-MF, 17%–21% on LMSYS) improve substantially over IT and RL but remain well above the backtracking methods, and emit incoherent text upon triggering.
@@
Table 2 extends this analysis to other adversarial strategies: Greedy Coordinate Gradient (GCG) attacks and manipulation of Decoding Parameters. These attacks represent different threat vectors, testing the model’s internal robustness and sensitivity to generation configurations. Against GCG attacks, RLBF consistently achieves the lowest ASR (4.3%–4.7%) compared to all baselines, including the strong Circuit Breakers (10.7%–13.4%) and BSAFE+ (5.7%–6.6%). Similarly, against Decoding Parameter attacks, while both BSAFE+ and RLBF perform exceptionally well (e.g., 1.0% ASR on MaliciousInstruct), RLBF shows a slight edge on the HEx-PHI benchmark (3.7% vs 5.0%). Neither the BSAFE+ data nor the RL stage contains GCG-style or decoding-exploit examples, so this robustness is obtained without training-time exposure to these attacks. Under the adaptive GCG variant of Zhang et al. (2025), which also optimizes the suffix to suppress backtracking tokens, RLBF’s ASR is unchanged (4.7% on AdvBench, 4.3% on HEx-PHI) and BSAFE+ shifts by at most 0.4 points. This superior performance against diverse, adaptive attacks further reinforces the benefits of the integrated RL optimization within RLBF, which likely fosters a more fundamental robustness to safety violations beyond what SFT-based correction or external filters alone can achieve.
@@ 5.4 Effect of Backtracking Capability in the Middle
5.5 Robustness to Critic Quality
To test whether the gains in Table 1 depend on a near-perfect critic, we repeated RL training with a weaker critic (LLaMA 3 8B with the same prompt, 82.0% accuracy on the validation set of Section 4.2.1). On LMSYS the resulting RLBF models reach 2%–4% ASR, compared to 1%–2% with the stronger critic and 14%–17% for BSAFE+. On LMSYS-MF the two critics yield comparable ASR (3%–7%), underscoring that the gains stem from policy learning with backtracking rather than critic accuracy alone.
5.6 Token Efficiency of Backtracking
We compared the “backtrack by $X$ tokens” command against BSAFE’s repeat-and-edit format. With BSAFE-style backtracking on identical data, BSAFE reaches 17%–22% ASR on LMSYS, between RL and BSAFE+. The data construction in Section 4.1.2 thus helps independently of the backtracking format. On responses where backtracking is triggered, BSAFE+ and RLBF save on average 68.3 and 72.6 tokens over BSAFE (74.8 and 79.1 tokens when the same corrections are replayed in BSAFE’s format). At typical serving speeds this corresponds to roughly 2.5–3.5 s saved per correction on hosted models and 5–10 s on edge devices, which is significant for interactive or voice-based deployments.